OpenStudy (anonymous):

This question has no label 1. Which system has (-8, 0) as the solution? 2. Which system has (1, 3) as the solution? 3. Which system has (-2, -2) as the solution? 4. Which system has no solutions? 5. Which system has infinite solutions? a. x - y = -2 3x - y = 0 b. x + y = -1 4x + 4y = -4 c. 4x + 6y = -12 2x + 3y = 9 d. x - 4y = -8 2x - 3y = -16 e. x + y = -4 -3x + 2y = 2

OpenStudy (anonymous):

this is a matching problem can aNYone help

OpenStudy (texaschic101):

(a) x - y = -2 --> x = y - 2 3x - y = 0 now sub y - 2 in for x 3(y - 2) - y = 0 3y - 6 - y = 0 2y = 6 y = 3 now sub 3 in for y x - y = -2 x - 3 = -2 x = -2 + 3 x = 1 answer is (1,3) which is # 2 do you understand how I did this or do you have questions ?

OpenStudy (texaschic101):

(b) x + y = -1 -->x = -y -1 4x + 4y = -4 now sub -y -1 in for x 4(-y-1) + 4y = -4 -4y - 4 + 4y = -4 -4 = -4 Answer is infinite solutions which is # 5

OpenStudy (texaschic101):

well...there is 2 of them for you....what you need to do is try to isolate a variable in one of your equations, then sub that back into the other equation and solve for the variable. Once one variable is found, sub that into the equation to find the other variable. Use the problems I did as your examples.

OpenStudy (anonymous):

so how would i figue them out

OpenStudy (anonymous):

so would my answers look like this 1=d 2=a 3=c 4=e 5=b

OpenStudy (mathstudent55):

1 & 2 are correct. 3 is not c. 4 is not e. 5 is correct.